# Numerical study of hypershadows in higher–dimensional black holes

### v2

@@ Abstract
We develop a fully numerical framework to compute and visualize the hypershadow[Novo:2024wyn], the three–dimensional generalization of the black hole shadow in five–dimensional spacetimes. Our method is based on backward ray tracing and allows flexible control over observer position, enabling the reconstruction of the full shadow volume. For visualization, we combine discrete sampling with surface contouring and introduce reflection difference maps on central slices to quantify mirror symmetries.
Applying this method to the Schwarzschild–Tangherlini and Myers–Perry geometries, we validate the former’s spherical symmetry and systematically discuss the hypershadow’s dependence on observer position and black hole spin parameters. We also provide compact quantitative measures for size reduction and global displacement, revealing clear monotonic trends. The framework is readily extendible to other metrics and opens the way to numerical studies of more exotic objects, such as black rings and their prospective toroidal hypershadows.
@@ 1 Introduction
In recent years, there has been growing interest in the properties of gravity beyond four dimensions, driven largely by developments in string theory, the AdS/CFT correspondence, and models with TeV–scale extra dimensions[strominger1996microscopic, aharony2000large, cavaglia2003black, kanti2004black]. In higher–dimensional spacetimes, black hole solutions exhibit qualitatively new features: the constraint of spherical horizon topology is relaxed, and the classical uniqueness theorems of four–dimensional general relativity no longer apply. Several higher–dimensional solutions to Einstein’s equations are well known, including the Tangherlini generalizations [Tangherlini:1963bw] of the Schwarzschild and Reissner–Nordström black holes for $D>4$ and the Myers–Perry extension [Myers:1986un] of the Kerr solution. All of these have spherical event horizons. However, they can coexist with solutions like the black ring [Emparan:2008eg, emparan2002rotating], which possesses a toroidal $S^{1}\times S^{D-3}$ topology, yet shares the same mass and angular momentum–providing a direct violation of uniqueness in vacuum gravity.
@@
As this example illustrates, adding just one extra dimension can fundamentally alter the solution space of black holes. This motivates a closer examination of which structural and observable features of black holes persist or fail to persist in higher dimensions. One particularly intriguing aspect is the black hole shadow, whose properties in higher dimensions remain only partially understood. The black hole shadow is the apparent image of a black hole projected onto the observer’s celestial sphere. In a four–dimensional spacetime, the shadow is a two–dimensional region on the surface of the celestial sphere. In a five–dimensional spacetime, the black hole shadow lies on a three–dimensional celestial sphere, giving rise to a genuinely three–dimensional structure, which we refer to as the hypershadow.
@@
There are two common approaches to computing black hole shadows: analytical and numerical. The analytical method typically uses the Hamilton–Jacobi formalism to derive separable geodesic equations and compute the shadow contour. The numerical method relies on backward ray tracing to reconstruct the shadow shape by simulating photon trajectories from the observer’s screen. While both methods have been extensively applied to two–dimensional shadows in four–dimensional spacetimes[bardeen1973timelike, cunha2018shadows, Papnoi:2014aaa], the three–dimensional hypershadow has so far only been explored analytically, most notably in Ref. [Novo:2024wyn, estrada2025five]. A fully numerical approach to constructing hypershadows remains undeveloped. In this work, we introduce such a numerical method to compute the shape of the hypershadow in higher–dimensional spacetimes.
@@
With the numerical method developed here, we are able to compute hypershadows for a broader class of black hole spacetimes. This allows us to systematically study how hypershadows deform in response to different observer position, such as changes in orientation, shape and position.
@@
To enhance visual interpretation, we propose an improved visualization scheme that combines discrete point sampling with surface contouring, offering clearer representations of the three–dimensional structure of hypershadows. Furthermore, by analyzing the reflection difference of their central slices, we introduce a quantitative measure of mirror symmetry in hypershadows. We then discuss the effects of observer position on the black hole shadow. These methodological advances lay the groundwork for future investigations into more exotic black hole geometries. In particular, they open the possibility of numerically exploring whether the hypershadow of a black ring exhibits a topological torus structure.
@@
A key novelty of this work lies in the systematic analysis of the observer inclination angle $\theta_{\rm o}$ in higher dimensions. In five–dimensional Myers–Perry black holes we introduce quantitative measures of hypershadow deformation: a distortion parameter $\delta_{s}$, capturing size reduction relative to the Schwarzschild–Tangherlini baseline, and a displacement parameter $\eta$, relevant in the singly rotating case. We find that $\delta_{s}$ grows with the spin and, in the single-rotation geometry, also with inclination, while $\eta$ quantifies a symmetry-breaking shift absent in the Cohomogeneity–one case. Together, these parameters show how $\theta_{\rm o}$ influences the rotational structure of higher–dimensional black holes.
@@
This paper is organized as follows. In Section 2, we introduce the backward ray-tracing method used to compute the hypershadow. In Section 3, we present the numerical hypershadows of several higher–dimensional black hole solutions. In Section 4, we study the effects of observer position on the black hole shadow. In Section 5, We present our conclusions.
2 Backward Ray-tracing
@@ 2.1 Local observer basis
The notation for the coordinate set $\chi=\{t,x,\theta,\phi,\psi\}$ was chosen in order to appear spherical-like. The radial coordinate $x$ is the square of the usual spherical coordinate $r$, i.e. $x=r^{2}$. This will allow us to simplify calculations and make many of the expressions more compact. The time coordinate $t$ and the azimuthal coordinates $\phi$ and $\psi$ adapt to the corresponding Killing vector, namely $\partial_{t},\partial_{\phi}$ and $\partial_{\psi}$. There is a conserved quantity of motion along geodesics associated with each Killing vector, $p_{t}=-E$, $p_{\phi}=\Phi$ and $p_{\psi}=\Psi$, respectively.
In this work, we adopt a particular orthonormal basis associated with a reference frame that has zero axial angular momentum with respect to spatial infinity. Such a frame is commonly referred to as the ZAMO (Zero Angular Momentum Observer) frame[Bardeen:1973tla], and is frequently used in black hole spacetimes to describe locally non–rotating observers. The ZAMO frame defines a natural decomposition of the observer’s orthonormal basis vectors in terms of the coordinate basis vectors $\{\partial_{t},\partial_{x},\partial_{\theta},\partial_{\phi},\partial_{\psi}\}$. While this decomposition is not unique, we adopt the following specific form in this paper:
@@
It is always possible to choose the non–Killing coordinates to be orthogonal so the corresponding coefficients would read:
@@
The locally measured momenta are:
@@ 2.2 Impact parameters
The shadow of a 4–dimensional black hole is typically represented as a two–dimensional shape on an image plane with Cartesian coordinates $\{X,Y\}$, which correspond to two observation angles $\{\alpha,\beta\}$. In the 5–dimensional case, this generalizes to a three-dimensional hypershadow, represented as a volume in Cartesian coordinates $\{X,Y,Z\}$, where the additional coordinate $Z$ corresponds to a third observation angle $\zeta$.
We therefore construct a 3D image space using Cartesian coordinates $\{X,Y,Z\}$, which represent the impact parameters of light rays and are proportional to the observation angles $\{\alpha,\beta,\zeta\}$. The observer is assumed to be located at coordinates $(x_{\rm o},\theta_{\rm o})$, positioned very far from the black hole, i.e. $x_{\rm o}\gg\mu$, where $\mu$ denotes the mass scale of the black hole. Throughout the paper we refer to $x_{\rm o}$ as the observer radial coordinate and to $\theta_{\rm o}$ as the observer inclination angle; their influence on the hypershadow will be analyzed in Sec. 4.
In the far–away observation limit $x_{\rm o}\to\infty$, the observation angles decay as $\sim 1/r_{\rm o}=1/\sqrt{x_{\rm o}}$, following the same scaling law as in the four–dimensional case. Accordingly, the impact parameters are defined as:
@@
Therefore, the spatial components of the photon’s momentum can be parameterized in terms of the observation angles as follows:
@@
Here, the origin of the observation angles, $\{\alpha,\beta,\zeta\}=\{0,0,0\}$, corresponds to a direction pointing radially away from the black hole. In the far–distance limit, the observation angles become very small, and the spatial momentum components can be approximated as:
@@
In natural units, the local energy of the photon is measured as $\varepsilon=\mathbf{p}$, which is also equal to the time component of the four–momentum, i.e. $\varepsilon=p^{(t)}$. Therefore, the impact parameters can be written as:
@@ 2.3 Geodesic Equation and Initial Condition
Since $p_{t}$, $p_{\phi}$, and $p_{\psi}$ are conserved quantities, they can be identified with E-E, $\Phi$, and $\Psi$, respectively. This leads to a system of seven first order differential equations in terms of the canonical variables $\{\chi^{\mu},p_{\mu}\}$. Such a formulation is particularly suitable for numerical geodesic integration, which determines whether a given light ray trajectory intersects the black hole.
@@
Our focus is thus on determining the initial conditions for the non–conserved momentum components $p_{x}(0)$ and $p_{\theta}(0)$. By combining the 5–momentum projections Eqs. (2.8) with the spatial momentum norm given in Eq. (2.7), we obtain:
@@
In traditional two–dimensional black hole shadow visualizations, such as those in Kerr spacetime[Cunha:2016bpi], the image is constructed by dividing the observation plane into small square pixels, each colored according to whether the corresponding light ray falls into the black hole. By increasing the resolution (i.e. using more, smaller pixels), one can obtain a sharp and detailed shadow image. This method works well in 2D, where the shadow is confined to a plane.
@@
Figure 1: Two–dimensional Kerr black hole shadows visualized with different pixel resolutions.
Fig. 1(a) uses a lower resolution, while Fig. 1(b) employs a higher resolution, resulting in a sharper and more refined shadow boundary. Here, $N_{\text{pix}}$ denotes the number of pixels along each spatial axis of the image grid. This illustrates the effectiveness of traditional 2D shadow visualization techniques when the shadow is confined to a plane.
However, when this approach is extended to visualize three-dimensional hypershadows in higher–dimensional spacetimes, the pixel–based method becomes inadequate. Even if we assign transparent colors to points where light rays escape and reserve opaque (e.g., black) colors for infalling ones, the result is still a solid, opaque volume of black voxels. This obscures the structure of the hypershadow, making it impossible to intuitively perceive its geometric shape in three dimensions.
To overcome this limitation, we propose a point-based visualization method. Instead of coloring volumetric pixels, we represent each sampling point as a discrete dot. The black dots correspond to light rays that fall into the black hole, while transparent points indicate escape. To enhance visibility of the hypershadow’s shape, we further connect the boundary of the black region using blue line segments. This approach reveals not only the outer boundary but also the overall spatial structure of the hypershadow, offering a clearer and more informative visualization than traditional voxel-based rendering.
@@ 3 Hypershadow
In this section, we present the numerical hypershadows corresponding to several five-dimensional black hole solutions, including the Schwarzschild–Tangherlini and Myers–Perry geometries. For each case, we reconstruct the three-dimensional shadow using the ray-tracing method developed earlier, and examine its global structure and symmetry properties. The aim is to demonstrate the effectiveness of our method and to highlight the distinct hypershadow features associated with static and rotating spacetimes.
3.1 Schwarzschild–Tangherlini Black hole
In 1963, Tangherlini derived this static, spherically symmetric and asymptotically flat vacuum solution as a generalization of the Schwarzschild black hole [Tangherlini:1963bw]. Strong gravitational lensing in the Tangherlini spacetime has been analyzed in  [Tsukamoto:2014dta, Singh:2017vfr]. We first introduce the five–dimensional Schwarzschild–Tangherlini black hole:
@@ 3.1 Schwarzschild–Tangherlini Black hole
Throughout this work we set $\mu=1$ for convenience, and the black hole horizon $x_{H}$ is obtained by solving $g_{tt}(x_{H})=0$, giving $x_{H}=\mu^{2}$.
@@
Figure 2: Hypershadows of the Schwarzschild–Tangherlini black hole rendered with different $N_{\rm pix}$. The observer position (the initial condition) is located at $\{65,\pi/2-1/20\}$.
As shown in Fig. 2, the Schwarzschild–Tangherlini black hole hypershadow appears perfectly spherical from the three–dimensional perspective. To rigorously confirm this spherical symmetry, we fix the observation angles to $\{\alpha,\beta,\zeta\}=0$ and compute the cross–sectional slices through the center of the hypershadow in the $XY$, XZXZ, and $YZ$ planes.
@@
[…]
For each slice we reflect the image about the indicated axis, then take the absolute difference with the original to highlight departures from exact mirror symmetry. Black pixels mark locations where the reflected slice differs from the original hypershadow slice. For the $XZ$ and $YZ$ slices the same convention applies, i.e. horizontal means reflection across the vertical axis of the displayed slice, and vertical means across the horizontal axis.
[…]
These findings not only confirm the expected spherical symmetry of the Schwarzschild–Tangherlini hypershadow, but also validate the precision and reliability of our numerical ray-tracing framework. This benchmark result serves as a foundation for exploring more intricate geometries in higher–dimensional rotating spacetimes.

### v3

# Numerical study of hypershadows in higher-dimensional black holes

@@ Abstract
We develop a fully numerical framework to compute and visualize the hypershadow[18], the three-dimensional generalization of the black hole shadow in five-dimensional spacetimes. Our method is based on backward ray tracing and allows flexible control over observer position, enabling the reconstruction of the full shadow volume. For visualization, we combine discrete sampling with surface contouring and introduce reflection difference maps on central slices to quantify mirror symmetries.
Applying this method to the Schwarzschild-Tangherlini and Myers-Perry geometries, we validate the former’s spherical symmetry and systematically discuss the hypershadow’s dependence on observer position and black hole spin parameters. We also provide compact quantitative measures for size reduction and global displacement, revealing clear monotonic trends. The framework is readily extendible to other metrics and opens the way to numerical studies of more exotic objects, such as black rings and their prospective toroidal hypershadows.
@@ 1 INTRODUCTION
In recent years, there has been growing interest in the properties of gravity beyond four dimensions, driven largely by developments in string theory, the AdS/CFT correspondence, and models with TeV–scale extra dimensions[21, 1, 5, 15]. In higher-dimensional spacetimes, black hole solutions exhibit qualitatively new features: the constraint of spherical horizon topology is relaxed, and the classical uniqueness theorems of four-dimensional general relativity no longer apply. Several higher-dimensional solutions to Einstein’s equations are well known, including the Tangherlini generalizations [22] of the Schwarzschild and Reissner-Nordström black holes for $D>4$ and the Myers-Perry extension [16] of the Kerr solution. All of these have spherical event horizons. However, they can coexist with solutions like the black ring [9, 10], which possesses a toroidal $S^{1}\times S^{D-3}$ topology yet shares the same mass and angular momentum—providing a direct violation of uniqueness in vacuum gravity.
@@
As this example illustrates, adding just one extra dimension can fundamentally alter the solution space of black holes. This motivates a closer examination of which structural and observable features of black holes persist or fail to persist in higher dimensions. One particularly intriguing aspect is the black hole shadow, whose properties in higher dimensions remain only partially understood. The black hole shadow is the apparent image of a black hole projected onto the observer’s celestial sphere. In a four-dimensional spacetime, the shadow is a two-dimensional region on the surface of the celestial sphere. In a five-dimensional spacetime, the black hole shadow lies on a three-dimensional celestial sphere, giving rise to a genuinely three-dimensional structure, which we refer to as the hypershadow.
@@
There are two common approaches to computing black hole shadows: analytical and numerical. The analytical method typically uses the Hamilton-Jacobi formalism to derive separable geodesic equations and compute the shadow contour. The numerical method relies on backward ray tracing to reconstruct the shadow shape by simulating photon trajectories from the observer’s screen. While both methods have been extensively applied to two-dimensional shadows in four-dimensional spacetimes[3, 7, 19, 24], the three-dimensional hypershadow has so far only been explored analytically, most notably in Refs. [18, 11]. A fully numerical approach to constructing hypershadows remains undeveloped. In this work, we introduce such a numerical method to compute the shape of the hypershadow in higher-dimensional spacetimes.
@@
With the numerical method developed here, we are able to compute hypershadows for a broader class of black hole spacetimes. This allows us to systematically study how hypershadows deform in response to different observer position, such as changes in orientation, shape, and position.
@@
To enhance visual interpretation, we propose an improved visualization scheme that combines discrete point sampling with surface contouring, offering clearer representations of the three-dimensional structure of hypershadows. Furthermore, by analyzing the reflection difference of their central slices, we introduce a quantitative measure of mirror symmetry in hypershadows. We then discuss the effects of observer position on the black hole shadow. These methodological advances lay the groundwork for future investigations into more exotic black hole geometries. In particular, they open the possibility of numerically exploring whether the hypershadow of a black ring exhibits a topological torus structure.
@@
A key novelty of this work lies in the systematic analysis of the observer inclination angle $\theta_{\rm o}$ in higher dimensions. In five-dimensional Myers-Perry black holes we introduce quantitative measures of hypershadow deformation: a distortion parameter $\delta_{s}$, capturing size reduction relative to the Schwarzschild-Tangherlini baseline, and a displacement parameter $\eta$, relevant in the singly rotating case. We find that $\delta_{s}$ grows with the spin and, in the single-rotation geometry, also with inclination, while $\eta$ quantifies a symmetry-breaking shift absent in the cohomogeneity-one case. Together, these parameters show how $\theta_{\rm o}$ influences the rotational structure of higher-dimensional black holes.
@@
This paper is organized as follows. In Sec. 2, we introduce the backward ray-tracing method used to compute the hypershadow. In Sec. 3, we present the numerical hypershadows of several higher–dimensional black hole solutions. In Sec. 4, we study the effects of observer position on the black hole shadow. In Sec. 5, we present our conclusions.
2 BACKWARD RAY TRACING
@@ 2.1 Local observer basis
The notation for the coordinate set $\chi=\{t,x,\theta,\phi,\psi\}$ was chosen in order to appear spherical-like. The radial coordinate $x$ is the square of the usual spherical coordinate $r$, i.e. $x=r^{2}$. This will allow us to simplify calculations and make many of the expressions more compact. The time coordinate $t$ and the azimuthal coordinates $\phi$ and $\psi$ adapt to the corresponding Killing vector, namely, $\partial_{t},\partial_{\phi}$, and $\partial_{\psi}$. There is a conserved quantity of motion along geodesics associated with each Killing vector, $p_{t}=-E$, $p_{\phi}=\Phi$, and $p_{\psi}=\Psi$, respectively.
In this work, we adopt a particular orthonormal basis associated with a reference frame that has zero axial angular momentum with respect to spatial infinity. Such a frame is commonly referred to as the zero angular momentum observer(ZAMO) frame[2], and is frequently used in black hole spacetimes to describe locally nonrotating observers. The ZAMO frame defines a natural decomposition of the observer’s orthonormal basis vectors in terms of the coordinate basis vectors $\{\partial_{t},\partial_{x},\partial_{\theta},\partial_{\phi},\partial_{\psi}\}$. While this decomposition is not unique, we adopt the following specific form in this paper:
@@
It is always possible to choose the non-Killing coordinates to be orthogonal so the corresponding coefficients would read
@@
The locally measured momenta are
@@ 2.2 Impact parameters
The shadow of a four-dimensional black hole is typically represented as a two-dimensional shape on an image plane with Cartesian coordinates $\{X,Y\}$, which correspond to two observation angles $\{\alpha,\beta\}$. In the five-dimensional case, this generalizes to a three-dimensional hypershadow, represented as a volume in Cartesian coordinates $\{X,Y,Z\}$, where the additional coordinate $Z$ corresponds to a third observation angle $\zeta$.
We, therefore, construct a 3D image space using Cartesian coordinates $\{X,Y,Z\}$, which represent the impact parameters of light rays and are proportional to the observation angles $\{\alpha,\beta,\zeta\}$. The observer is assumed to be located at coordinates $(x_{\rm o},\theta_{\rm o})$, positioned very far from the black hole, i.e., $x_{\rm o}\gg\mu$, where $\mu$ denotes the mass scale of the black hole. Throughout the paper we refer to $x_{\rm o}$ as the observer radial coordinate and to $\theta_{\rm o}$ as the observer inclination angle; their influence on the hypershadow will be analyzed in Sec. 4.
In the far-away observation limit $x_{\rm o}\to\infty$, the observation angles decay as $\sim 1/r_{\rm o}=1/\sqrt{x_{\rm o}}$, following the same scaling law as in the four-dimensional case. Accordingly, the impact parameters are defined as
@@
Therefore, the spatial components of the photon’s momentum can be parametrized in terms of the observation angles as follows:
@@
Here, the origin of the observation angles, $\{\alpha,\beta,\zeta\}=\{0,0,0\}$, corresponds to a direction pointing radially away from the black hole. In the far–distance limit, the observation angles become very small, and the spatial momentum components can be approximated as
@@
In natural units, the local energy of the photon is measured as $\varepsilon=\mathbf{p}$, which is also equal to the time component of the four-momentum, i.e. , $\varepsilon=p^{(t)}$. Therefore, the impact parameters can be written as
@@ 2.3 Geodesic equation and initial condition
Since $p_{t}$, $p_{\phi}$, and $p_{\psi}$ are conserved quantities, they can be identified with E-E, $\Phi$, and $\Psi$, respectively. This leads to a system of seven first-order differential equations in terms of the canonical variables $\{\chi^{\mu},p_{\mu}\}$. Such a formulation is particularly suitable for numerical geodesic integration, which determines whether a given light ray trajectory intersects the black hole.
@@
Our focus is thus, on determining the initial conditions for the nonconserved momentum components $p_{x}(0)$ and $p_{\theta}(0)$. By combining the five-momentum projections Eqs. (2.8) with the spatial momentum norm given in Eq. (2.7), we obtain
@@
In traditional two-dimensional black hole shadow visualizations, such as those in Kerr spacetime[6], the image is constructed by dividing the observation plane into small square pixels, each colored according to whether the corresponding light ray falls into the black hole. By increasing the resolution (i.e. using more, smaller pixels), one can obtain a sharp and detailed shadow image. This method works well in 2D, where the shadow is confined to a plane.
@@
Figure 1: Two-dimensional Kerr black hole shadows visualized with different pixel resolutions.
Fig. (1(a)) uses a lower resolution, while Fig. (1(b)) employs a higher resolution, resulting in a sharper and more refined shadow boundary. Here, $N_{\text{pix}}$ denotes the number of pixels along each spatial axis of the image grid. This illustrates the effectiveness of traditional 2D shadow visualization techniques when the shadow is confined to a plane.
However, when this approach is extended to visualize three-dimensional hypershadows in higher-dimensional spacetimes, the pixel-based method becomes inadequate. Even if we assign transparent colors to points where light rays escape and reserve opaque (e.g., black) colors for infalling ones, the result is still a solid, opaque volume of black voxels. This obscures the structure of the hypershadow, making it impossible to intuitively perceive its geometric shape in three dimensions.
To overcome this limitation, we propose a point-based visualization method. Instead of coloring volumetric pixels, we represent each sampling point as a discrete dot. The black dots correspond to light rays that fall into the black hole, while transparent points indicate escape. To enhance visibility of the hypershadow’s shape, we further connect the boundary of the black region using blue line segments. This approach reveals not only the outer boundary, but also the overall spatial structure of the hypershadow, offering a clearer and more informative visualization than traditional voxel-based rendering.
@@ 3 HYPERSHADOW
In this section, we present the numerical hypershadows corresponding to several five-dimensional black hole solutions, including the Schwarzschild-Tangherlini and Myers-Perry geometries. For each case, we reconstruct the three-dimensional shadow using the ray-tracing method developed earlier and examine its global structure and symmetry properties. The aim is to demonstrate the effectiveness of our method and to highlight the distinct hypershadow features associated with static and rotating spacetimes.
3.1 Schwarzschild-Tangherlini black hole
In 1963, Tangherlini derived this static, spherically symmetric and asymptotically flat vacuum solution as a generalization of the Schwarzschild black hole [22]. Strong gravitational lensing in the Tangherlini spacetime has been analyzed in  [23, 20]. We first introduce the five-dimensional Schwarzschild-Tangherlini black hole:
@@ 3.1 Schwarzschild-Tangherlini black hole
Throughout this work, we set $\mu=1$ for convenience, and the black hole horizon $x_{H}$ is obtained by solving $g_{tt}(x_{H})=0$, giving $x_{H}=\mu^{2}$.
@@
Figure 2: Hypershadows of the Schwarzschild-Tangherlini black hole rendered with different $N_{\rm pix}$. The observer position (the initial condition) is located at $\{65,\pi/2-1/20\}$.
As shown in Fig. 2, the Schwarzschild-Tangherlini black hole hypershadow appears perfectly spherical from the three-dimensional perspective. To rigorously confirm this spherical symmetry, we fix the observation angles to $\{\alpha,\beta,\zeta\}=0$ and compute the cross-sectional slices through the center of the hypershadow in the $XY$, XZXZ, and $YZ$ planes.
@@
Figure 3: (a) Central slices in the $XY$, XZXZ, and $YZ$ planes. (b) H-$\Delta$: difference to the left-right mirror. (c) V-$\Delta$: difference to the up-down mirror. Resolution $N_{\rm pix}=512$.
The symbols “H-$\Delta$” and “V-$\Delta$” denote pixelwise difference maps under mirror reflections of each 2D slice. Specifically, for a slice plotted with the $X$ axis horizontal and $Y$ axis vertical:
  • horizontal reflection (H-$\Delta$) is the left-right flip across the vertical midline $X\!\mapsto\!-X$, and
  • vertical reflection (V-$\Delta$) is the up-down flip across the horizontal midline $Y\!\mapsto\!-Y$.
For each slice, we reflect the image about the indicated axis and then take the absolute difference with the original to highlight departures from exact mirror symmetry. Black pixels mark locations where the reflected slice differs from the original hypershadow slice. For the $XZ$ and $YZ$ slices, the same convention applies i.e., horizontal means reflection across the vertical axis of the displayed slice, and vertical means across the horizontal axis.
Figure. 3 displays the central slices together with H-$\Delta$ and V-$\Delta$ (definitions above). The vanishing of both maps (up to numerical tolerance) indicates exact reflection symmetry in each coordinate plane.
These findings not only confirm the expected spherical symmetry of the Schwarzschild-Tangherlini hypershadow, but also validate the precision and reliability of our numerical ray-tracing framework. This benchmark result serves as a foundation for exploring more intricate geometries in higher-dimensional rotating spacetimes.